OpenStudy (anonymous):

elimination method 0.3x-0.2y=4 0.2x+0.5y= -21/23

OpenStudy (anonymous):

multiply first equn. by 2 and second by 3 0.6x - 0.4y = 8.....................(1) 0.6x + 1.5y = -63/23..........(2) (2) - (1) :- 1.9y = -63/23 - 184/23 1.9y = -247/23 y = -5.652 substitute y = -5.652 into equation. 1: 0.6x - (0.4 X -5.652) = 8 x =[ 8 + (0.4 X -5.652] / 0.6 =9.5653